OpenStudy (anonymous):

find the solution of the square root of the quantity of x plus 3 plus 4 equals 6 and determine if it is an extraneous solution

OpenStudy (anonymous):

\[\sqrt{x+3}+4=6\] ???

OpenStudy (anonymous):

solution: 2, not extraneous

OpenStudy (perl):

Is this the question? $$\Large { \sqrt{x+3}+4=6 } $$

OpenStudy (anonymous):

yes @perl

OpenStudy (perl):

$$ \Large { \sqrt{x+3}+4=6 \\ \therefore \\ \sqrt{x+3}=6 - 4 \\ \therefore \\ \sqrt{x+3}=2 \\ \therefore \\ (\sqrt{x+3})^2=2^2 } $$

OpenStudy (anonymous):

that isnt one of the answer choices

OpenStudy (perl):

please finish the problem :)

OpenStudy (preetha):

Lisa, Perl is only trying to help you learn. Please try it out and he can help you some more.

OpenStudy (anonymous):

Idk how to do it which is why I am here

OpenStudy (perl):

did the last step make sense

OpenStudy (anonymous):

no its where I got lost

OpenStudy (perl):

ok does it make sense up to here, we can go back a step or two

OpenStudy (anonymous):

yes it does

OpenStudy (perl):

ok great :)

OpenStudy (perl):

$$ \Large { \sqrt{x+3}+4=6 \\ \therefore \\ \sqrt{x+3}=6 - 4 \\ \therefore \\ \sqrt{x+3}=2 \\ \therefore \\ (\sqrt{x+3})^2=2^2 \\ \therefore \\ x+3=2^2 } $$

OpenStudy (perl):

because \( \large (\sqrt{u})^2= u \)

OpenStudy (perl):

squaring a square root cancels , for example \( \Large (\sqrt{4}) ^2 = 4\)

OpenStudy (anonymous):

Alright so far I understand so the answer would be 1 not extraneous

OpenStudy (perl):

the last step is to test the solution in the original equation

OpenStudy (perl):

and that is correct,

OpenStudy (anonymous):

really wow thanks so much

OpenStudy (perl):

$$ \Large { \sqrt{\color{red}1+3}+4=6 } $$

OpenStudy (perl):

your welcome :)

OpenStudy (anonymous):

What do. You. Guys. Need help qith huh